OpenStudy (tayhay36):

Write a two column proof. Given: 7y=8x-14 and y=6 Prove: x=7 Please help me. I'm so bad at math.

OpenStudy (anonymous):

LS : 7 x 6 = 42 RS 8 x 7 - 14 = 56 - 14 = 42 42 = 42 LS = RS Therefore x, y = 7 , 6